# Neuromorphic Intermediate Representation: A Unified Instruction Set for Interoperable Brain-Inspired Computing

### v1

@@ 2.1.1 Leaky integrate-and-fire dynamics
2.1.2 Convolutional neural network
For the second experiment, we studied a 9-layer spiking convolutional neural network (SCNN) using 2-dimensional convolutions, sum pooling, and integrate-and-fire neurons, shown in Figure 3 (c). The accuracies of the SCNNs across the different platforms on the withheld test dataset are shown to be very similar, see Table 1, with a mean test accuracy of 98.1% and a standard deviation of only 0.1% across all platforms. These results demonstrate that NIR can effectively act as an intermediate representation of a trained, rate-based spiking convolutional neural network, as the end-task performance is preserved across all tested platforms.
We further analyze the activities of the SCNN’s hidden layers across all different platforms. For a visually interpretable comparison, we compare the time-averaged activities for the first channel of the first hidden layer of I&F neurons in Figure 4 (b). The similarity measure tells us that most platforms have almost identical dynamics, while particularly Rockpool, Sinabs, snnTorch, and Speck deviates. Since the accuracies are relatively close, we observe that the rate encoding and feed-forward architecture of the investigated SCNN are relatively robust to these mismatches.
@@ 2.1.2 Convolutional neural network
2.1.3 Recurrent neural network
@@
For the third and final task, we studied a two-layer spiking recurrent neural network (SRNN), with populations of current-based LIF (CuBa-LIF) neurons. The network was trained in snnTorch with backpropagation-through-time (BPTT) and surrogate gradients to perform a Braille letter recognition task [35]. After training, the network was exported to NIR, and evaluated on all other platforms. The accuracies and similarity measures for their respective activations are shown in Figure 4. We train two different networks, depending on the discretization of the reset mechanisms for post-spike membrane reset (shown in Figure 2 and equations (5) and (6)). Note that the NIR graph structure is the same, only the parameter values differ.
@@
In this experiment, the performance across different platforms was notably less consistent than in the prior two tasks. Contrarily, the SRNN presented here displayed significant disparities in end-task performance when implemented on different platforms. This divergence can be attributed to the recurrent connections within the network, which tend to accentuate subtle discrepancies in the dynamics. The SRNN’s training on snnTorch suggests that its dynamics are intricately calibrated to the specificities of this simulator. Therefore, deploying the network on a platform with dissimilarities from snnTorch’s nuances can introduce pronounced differences in the system’s dynamics, thereby affecting the end-task outcome.
@@
Hence, NIR emerges not only as a representation but also as an instrumental tool—it facilitates a deeper exploration of the discrepancies between platforms, revealing the influence of diverse parameterization, levels of precision, and models on the resultant performance mismatches. NIR thus offers a valuable avenue for research into the robustness of neural networks, specifically gauging their resilience to simulator and device mismatch, and thereby stands as a pivotal tool for the advancement of neuromorphic computing.
2.2 NIR axioms
@@
The nodes in NIR are described by hybrid continuous-time systems, capturing the evolution of some physical quantity as well as jumps and discrete resets. Rather than executing these dynamics, the NIR graph merely declares them, ensuring that any computational backend is equipped with complete information required to interpret and evaluate the represented computation. It is essential to clarify that NIR’s objective is not to rectify discrepancies arising from mismatches between platforms. Rather, NIR offers a common ground upon which all hardware platforms can address shared challenges and brings into sharp focus the inconsistencies that arise from a heterogeneous neuromorphic computing landscape where neuron models and discretization schemes are not standardized.
@@
    A computational node $v\in\mathcal{V}$ is composed of a primitive $c\in\mathcal{C}$ and its parameters $\theta$. This yields a complete description of an input-output transformation, and thus fully describes the computation implemented by the computational node $v$. For example, the computational node for an affine connection layer consists of the primitive $c_{\text{Affine}}$ along with the weight matrix $W$ and the bias $b$ as the parameters $\theta=(W,b)$, such that $v_{\text{Affine}}=(c_{\text{Affine}},\theta)$.
@@
    Edges do not perform any computation, and are effectively the identity map $I:x\mapsto x$.
@@
    One input port can receive multiple incoming edges which are then, by convention, summed together elementwise. Similarly, one output port can send its information via different edges to various input ports. The information is then passed by values, not by reference, such that information can only flow along the direction of an edge.
  1. (2.3)
    Every NIR graph has exactly one input node and one output node. These nodes do not transform their input—they exist to disambiguate where the input and output of a graph is situated.
The above axioms closely follow conventional descriptions of directed signal-flow graphs. Figure 2 further illustrates the workings of NIR graphs, and Figure 3 shows three example NIR graphs that are used in the experiments from Section 2.1.
2.3 Computational primitives
@@
Input
Input shape
$\mathbb{N}^{N_{0}\cdots N_{n}}$
Output
Output shape
$\mathbb{N}^{N_{0}\cdots N_{0}}$
Affine
$W,b$
$\mathbb{R}^{N_{0}\cdots N_{n}}\times\mathbb{R}^{N_{1}\cdots N_{n}}$
$W\,i(\cdot)+b$
Convolution
$W$, Stride, Padding, Dilation, Groups, Bias
$\mathbb{R}^{C_{\text{out}}\times C_{\text{in}}\times N_{0}\cdots N_{n}}\times\mathbb{N}^{N_{0}\cdots N_{0}}\times\mathbb{N}^{N_{0}\cdots N_{0}}\times\mathbb{N}^{N_{0}\cdots N_{0}}\times\mathbb{N}\times\mathbb{R}^{N_{0}\cdots N_{0}}$
$f\star g$
Delay
$\tau$
$\mathbb{R}^{N}$
$i(t-\tau)$
Flatten
Input shape, Start dim., End dim.
$\mathbb{N}^{N_{0}\cdots N_{n}}\times\mathbb{N}\times\mathbb{N}$
Integrator
$R$
$\mathbb{R}^{N_{0}\cdots N_{n}}$
$\dot{v}=R\,i(\cdot)$
Leaky integrator (LI)
$\tau$, $R$, $v_{leak}$
$\mathbb{R}^{N_{0}\cdots N_{n}}\times\mathbb{R}^{N_{0}\cdots N_{n}}\times\mathbb{R}^{N_{0}\cdots N_{n}}$
$\tau\dot{v}=(v_{leak}-v)+R\,i(\cdot)$
Linear
$W$
$\mathbb{R}^{N_{0}\cdots N_{n}}$
$W\,i(\cdot)$
Scale
$s$
$\mathbb{R}^{N_{0}\cdots N_{n}}$
$s\,i(\cdot)$
Spike
$\theta_{\text{thr}}$
$\mathbb{R}^{N_{0}\cdots N_{n}}$
$\delta(i(\cdot)-\theta_{\text{thr}})$
Higher-order primitive Computation
@@
[…]
Table 2: The computational primitives, their parameters and computational models (top). The bottom listing defines higher-order primitives as compositions. Parameters are typed for an arbitrary number of dimensions, $n$. $i(\cdot)$ denotes input at a given point in time, where $t$ denotes a specific time, $C_{\text{out}}\times C_{\text{in}}$ denotes convolutional output and input channels, and $\delta$ denotes the Dirac delta function.
NIR defines 11 computational primitives and 3 higher-order primitives, listed in Table 2. We define common neuromorphic components like the linear map, leaky integrator, and spike threshold function, but we further included mathematical primitives such as the affine map and convolution. The input and output nodes serve to disambiguate the entries and exits of a graph.
The 11 primitives in NIR are “fundamental” in the sense that the backends implementing the primitives are required to approximate the computation of the idealized description as closely as possible within the limitations of the platform. Any given platform is not expected to implement the full specification. This is particularly true for functionally specialized hardware, where hardware restrictions render certain functional primitives impossible. We detail such restrictions in Section 4.3.1.
NIR does not restrict the implementation apart from stating the idealized descriptions, and it is expected that the exact computation of the primitives will vary across platforms. This is particularly important for digital systems and their choice of integration and discretization method. Given the heterogeneity of the present hardware landscape, this amount of freedom becomes a practical choice as well; any restrictions towards discretizing the hybrid ODEs will invariably lead to incompatibilities and numerical deviations across platforms.
@@
    The integrate-and-fire (IF) neuron, defined as an integrator composed with a threshold and membrane reset mechanism.
@@
    The leaky integrate-and-fire (LIF) neuron, defined as a leaky integrator composed with a threshold and membrane reset mechanism.
@@ 3 Discussion
We presented an intermediate representation for continuous-time neuromorphic systems, dubbed the Neuromorphic Intermediate Representation (NIR), in which we established a set of primitives that can be arbitrarily composed to create sophisticated computational structures. We demonstrated the application of NIR by evaluating three tasks across the supported neuromorphic platforms: a leaky integrate-and-fire model, a spiking convolutional neural network using integrate-and-fire neurons, and a spiking recurrent neural network using current-based leaky integrate-and-fire neurons. We found that NIR faithfully represents the idealized underlying computation and that the execution aligns well across the supported platforms for feed-forward dynamics, such as a single LIF neuron or a spiking convolutional network. For more complex and time-sensitive models, such as the recurrent network, we found that platform-specific constraints and discretization choices produced somewhat diverging activation patterns and accuracies.
The possibility of conversion between different software systems allows a user to leverage features of different libraries such as quantization aware training or particularly fast training speeds. This cross-compatibility paves the way for a versatile, shared SNN tooling infrastructure.
Because NIR represents an idealized model, we argue that NIR helps highlight differences between neuromorphic hardware and software that would be difficult to identify otherwise. The common representation in NIR therefore not only permits the translation of neuromorphic computational models between systems, but also simplifies platform-specific optimizations. While such optimizations may reduce the ability for the model to generalize, it is an important component if neuromorphic systems are to compete with conventional compute platforms.
One crucial implication that we do not discuss in this work is the connection to physical and analog systems. The primitives in NIR are effectively first-order integrators and can be composed to express higher-order primitives, as in the LIF and CuBa-LIF models. In turn, this can be used to express more complex partial differential equations and model problems like diffusion, fluid dynamics, and quantum mechanics. Effectively, our method generalizes to other kinds of hardware, including analog substrates.
Our main goal has been to address the lack of shared representations for neuromorphic computing. NIR effectively guarantees reproducibility across neuromorphic simulators and hardware. Similar to how the initial design of digital instruction sets propelled the advent of modern-day computing, we believe this to be a crucial step towards further development of neuromorphic computing. We hope that NIR can accelerate both the academic development required to propel neuromorphic computing to perform on-par with present-day deep learning, while solving practical problems in the near future.
NIR is open-source, licensed under BSD-3, and available at https://github.com/neuromorphs/NIR

### v2

@@ 2 Results
Convolutional neural network
For the second experiment, we studied a 9-layer spiking convolutional neural network (SCNN) using 2-dimensional convolutions, sum pooling, and integrate-and-fire neurons, shown in Figure 3 (c). The SCNN was trained on the Neuromorphic MNIST (N-MNIST) dataset to recognize digits from 0 to 9 [70]. Details on the training and exact network architecture are shown in Section Spiking convolutional neural network. In our experiment, we compare not only the end-task accuracy, but also the activity of the first hidden layer.
The accuracies of the SCNNs across the different platforms on the withheld test dataset are similar (see Table 1), with a mean test accuracy of 97.7% and a standard deviation of 0.9% across all platforms. These results demonstrate that NIR can effectively act as an intermediate representation of a trained, rate-based spiking convolutional neural network, as the end-task performance is preserved across all tested platforms.
We further analyze the activities of the SCNN’s hidden layers across all different platforms. For a visually interpretable comparison, we use the cosine similarity of time-averaged activities, i.e. spike rates, for the first channel of the first hidden layer of I&F neurons in Figure 4 (b): $S_{c}(\mathbf{r}_{1},\mathbf{r}_{2})=(\mathbf{r}_{1}\cdot\mathbf{r}_{2})/(\|\mathbf{r}_{1}\|_{2}\|\mathbf{r}_{2}\|_{2})$ where $\mathbf{r}_{i}$ is the flattened vector of spike rates from platform $i$. The similarity measure tells us that most platforms have almost identical dynamics, while particularly Rockpool, Sinabs, snnTorch, and Speck deviates. Since the accuracies are relatively close, we observe that the rate encoding and feed-forward architecture of the investigated SCNN are relatively robust to these mismatches.
@@
Recurrent neural network
@@
For the third and final experiment, we studied a spiking recurrent neural network (SRNN) with one hidden layer, with populations of current-based LIF (CuBa-LIF) neurons. The network was trained in snnTorch with backpropagation-through-time (BPTT) and surrogate gradients to perform a Braille letter recognition task [71]. Further details are reported in Section Spiking recurrent neural network. After training, the network was exported to NIR, and evaluated on all other platforms. The accuracies and similarity measures for their respective activations are shown in Figure 4 (c-d). We trained two networks with different choices for the discretization of the reset mechanisms for post-spike membrane reset (shown in equations (1) and (2) in the Supplementary Material). Similar to the SCNN experiment, we measured not only the end-task accuracy, but also the activity of the hidden recurrent layer.
@@
In this experiment, the performance across different platforms was notably less consistent than in the prior two. Contrarily, the SRNN presented here displayed significant disparities in end-task performance when implemented on different platforms. This divergence can be attributed to the recurrent connections within the network, which tend to accentuate subtle discrepancies in the dynamics. The SRNN’s training in snnTorch suggests that its dynamics is intricately calibrated to the specifics of this simulator. Therefore, deploying the network on a platform with dissimilarities from snnTorch’s nuances can introduce pronounced differences in the system’s dynamics, thereby affecting the end-task outcome.
@@
In sum, NIR emerges not only as a representation but also as a measurement tool—it facilitates a deeper exploration of the discrepancies between platforms, revealing the influence of diverse parameterization, levels of precision, and models on the resultant performance mismatches. NIR thus offers a valuable avenue for further research by quantifying the robustness of neural networks, specifically gauging their resilience to simulator and device mismatch, and thereby stands as a pivotal tool for the advancement of neuromorphic computing.
Discussion of mismatches
Instead of providing perfect functional reproducibility, NIR establishes an idealized model that serves as a common ground for comparative analysis while highlighting inter-platform discrepancies. We attribute these discrepancies to three main causes:
Neuron model implementation. As shown in Section A in the Supplementary Material, the simulators and chips do not use the same neuron model definitions. Whereas some definitions can be made equal by matching parameters, differences in spike and reset timing or different discretization choices lead to changes in the neuron dynamics.
Quantization. The neuromorphic chips we tested use quantization to reduce on-chip memory requirements and power required for computation. Quantization will change the activity due to inherent rounding errors, which can cause performance degradation [72]. We only used post-training quantization to reveal discrepancies between the frameworks. Although quantization-aware training (QAT) or fine-tuning may be able to recover lost performance, we do not provide a detailed study of the extent to which QAT reduces mismatches here.
Determinism. Whereas simulation frameworks are usually deterministic in their computation, asynchronous hardware, such as Speck and SpiNNaker2, are not fully deterministic. This variability can lead to further activity discrepancies across platforms. Most event-based processing systems do not guarantee the determinism of the individual events.
NIR axioms
@@
The nodes in NIR are described by hybrid continuous-time systems, capturing the realistic evolution of some physical quantity as well as jumps and discrete resets. Instead of executing these dynamics, the NIR graph merely declares them, ensuring that any computational backend is equipped with complete information required to interpret and evaluate the represented computation. It is essential to clarify that NIR’s objective is not to rectify discrepancies arising from mismatches between platforms. Rather, NIR offers a common framework on which all hardware platforms can address shared challenges and brings into sharp focus the inconsistencies that arise from a heterogeneous neuromorphic computing landscape where neuron models and discretization schemes are not standardized.
@@
    A computational node $v\in\mathcal{V}$ is composed of a primitive $c\in\mathcal{C}$ and its parameters $\theta$. This yields a complete description of an input-output transformation, and thus fully describes the computation implemented by the computational node $v$. For example, the computational node representing an affine linear map consists of the primitive $c_{\text{Affine}}$ along with the weight matrix $W$ and the bias $b$ as the parameters $\theta=(W,b)$, such that $v_{\text{Affine}}=(c_{\text{Affine}},\theta)$.
@@
    Edges do not perform any computation and are effectively the identity map $I:x\mapsto x$.
@@
    One input port can receive multiple incoming edges which are then, by convention, summed together element-wise. Similarly, one output port can send its information via different edges to various input ports. The information is then passed by values, not by reference, such that information can only flow along the direction of an edge.
The above axioms closely follow conventional descriptions of directed signal-flow graphs. Figure 2 further illustrates the workings of NIR graphs, and Figure 3 shows three example NIR graphs that are used in the experiments from Section Experiments.
Computational primitives
@@
Input Input shape $\mathbb{N}^{N_{0}\cdots N_{n}}$
Output Output shape $\mathbb{N}^{N_{0}\cdots N_{0}}$
Affine $W,b$ $\mathbb{R}^{N_{0}\cdots N_{n}}\times\mathbb{R}^{N_{1}\cdots N_{n}}$ $W\,i(\cdot)+b$
Convolution $W$, Stride, Padding, Dilation, Groups, Bias $\mathbb{R}^{C_{\text{out}}\times C_{\text{in}}\times N_{0}\cdots N_{n}}\times\mathbb{N}^{N_{0}\cdots N_{0}}\times\mathbb{N}^{N_{0}\cdots N_{0}}\times\mathbb{N}^{N_{0}\cdots N_{0}}\times\mathbb{N}\times\mathbb{R}^{N_{0}\cdots N_{0}}$ $f\star g$
Delay $\tau$ $\mathbb{R}^{N}$ $i(t-\tau)$
Flatten Input shape, Start dim., End dim. $\mathbb{N}^{N_{0}\cdots N_{n}}\times\mathbb{N}\times\mathbb{N}$
Integrator $R$ $\mathbb{R}^{N_{0}\cdots N_{n}}$ $\dot{v}=R\,i(\cdot)$
Leaky integrator (LI) $\tau$, $R$, $v_{leak}$ $\mathbb{R}^{N_{0}\cdots N_{n}}\times\mathbb{R}^{N_{0}\cdots N_{n}}\times\mathbb{R}^{N_{0}\cdots N_{n}}$ $\tau\dot{v}=(v_{leak}-v)+R\,i(\cdot)$
Linear $W$ $\mathbb{R}^{N_{0}\cdots N_{n}}$ $W\,i(\cdot)$
Scale $s$ $\mathbb{R}^{N_{0}\cdots N_{n}}$ $s\,i(\cdot)$
Spike $\theta_{\text{thr}}$ $\mathbb{R}^{N_{0}\cdots N_{n}}$ $\delta(i(\cdot)-\theta_{\text{thr}})$
Higher-order primitive Composition
Integrate-and-fire (I&F) (Leaky) IntegratorSpikeLinear
Leaky integrate-and-fire (LIF)
@@
Table 2: Computational primitives in NIR. Top: The computational primitives, their parameters and computational models. Bottom: Higher-order primitives as compositions. Parameters are typed for an arbitrary number of dimensions, $n$. $i(\cdot)$ denotes input at a given point in time, where $t$ denotes a specific time, $C_{\text{out}}\times C_{\text{in}}$ denotes convolutional output and input channels, and $\delta$ denotes the Dirac delta function.
NIR defines 11 computational primitives and 3 higher-order primitives, listed in Table 2, where each primitive describes the evolution of a hybrid system over time (see Section NIR axioms). We define common neuromorphic components like the linear map, leaky integrator, and spike threshold function, but we further included mathematical primitives such as the affine map and convolution. The input and output nodes serve to disambiguate the entries and exits of a graph.
The 11 primitives in NIR are “fundamental” in the sense that the backends implementing the primitives are required to approximate the computation of the idealized description as closely as possible within the limitations of the platform. Any given platform is not expected to implement the full specification. This is particularly true for functionally specialized hardware, where hardware restrictions render certain functional primitives impossible. We detail such restrictions in Section Hardware platforms and how NIR can deal with hardware constraints in Modeling hardware constraints with NIR.
NIR does not restrict the implementation apart from stating the idealized descriptions, and it is expected that the exact computation of the primitives will vary across platforms. This is particularly important for digital systems and their choice of integration and discretization methods. Given the heterogeneity of the present hardware landscape, this amount of freedom becomes a practical choice as well; imposing any restriction on the integration scheme for the hybrid ODEs will invariably lead to incompatibilities and numerical deviations across platforms.
@@
    The integrate-and-fire (IF) neuron, defined as an integrator composed with a spike function and a feedback mechanism that acts as membrane reset.
@@
    The leaky integrate-and-fire (LIF) neuron, defined as a leaky integrator composed with a spike function and a feedback mechanism that acts as membrane reset.
@@
Relation between NIR and existing intermediate representations
The field of neuromorphic computing is divided into separate paths for digital and analog hardware developments, which often progress in isolation. By providing a unified framework that accommodates both of the paths, NIR transcends this divide, capturing the underlying essence of neuromorphic computing that is rooted in the use of time as a computational element and the exploitation of continuous-time dynamics where “time represents itself” [73]. To address these aspects, NIR draws inspiration from, and share commonalities with, previous efforts. Below, we sketch the unique characteristics of NIR in relation to other intermediate representations related to neuromorphic computing. A tabular overview is available in the Supplementary Material, Section B.
Modular approach to primitives: NIR stands out by not prescribing a single representation for computations, but instead offers a modular approach where primitives can be composed arbitrarily. The concept of composing primitives has proven successful in other frameworks like Fugu and Lava for the neuromorphic domain, and Compiler infrastructure projects like LLVM [60] and MLIR [15] for the conventional domain. Both LLVM and MLIR are designed to optimize discrete machine-level instructions, rather than continuous-time systems equations. Modularity and flexibility allow NIR graphs to be tailored to diverse hardware platforms, ensuring that they can be adapted to the constraints of each platform (see Section Modeling hardware constraints with NIR). Moreover, NIR expands the application of neuromorphic computing beyond neuroscience to include deep learning operations, enabling the modeling of both artificial neural networks (ANNs) and spiking neural networks (SNNs), as well as supporting hybrid ANN-SNN computations.
Agnostic to hardware/software platforms: NIR adopts a purely declarative style, making it agnostic to the specifics of any hardware or software platform. Any given backend can choose to interpret and implement NIR graphs, as long as the backend approximates NIR’s underlying continuous-time dynamics. This contrasts efforts like PyNN, NeuroML, and Fugu, which make assumptions about the runtime environments. NIR is designed to integrate with existing standards and compilers. This interoperability aligns NIR with model-centric APIs in deep learning and frameworks like MMdnn [74] or Ivy [18], facilitating the transfer of models and computations across different libraries and platforms.
Serializable exchange format: Currently, NIR functions primarily as a representation and exchange format, focusing on the portability of core model dynamics rather than optimizing hardware efficiency and model performance. This approach positions NIR similarly to ONNX [14] or MMdnn [74], contrasting with frameworks like MLIR [15] and TVM [17] that provide cross-compilation and multi-layer abstractions. As NIR’s base representations stabilize, there is potential to introduce additional layers of abstraction, enhancing its capability to serve as a foundational building block for future computational frameworks.
Applicability to multiple time scales: NIR and NeuroML model continuous-time systems, which translate trivially to continuous-time simulators, as shown in Figure 1 (a). By supporting different time constants in the same computational graph, the model can detail dynamics at multiple time scales simultaneously. Behaviors at multiple levels are challenging to capture with discrete simulators that advance via fixed time steps, particularly if they are complex or chaotic.
@@ 3 Discussion
We presented an intermediate representation for digital neuromorphic systems, the Neuromorphic Intermediate Representation (NIR), as a platform-independent set of continuous-time computational primitives. NIR can be viewed as a “neuromorphic instruction set” that can be composed arbitrarily. It is an ideal serializable representation for multiple neuromorphic platforms, serving as a storage format for software platforms and a source language for these hardware systems. We documented the support of NIR across 11 different neuromorphic platforms on which we demonstrated three specific computational graphs: a leaky integrate-and-fire neuron, a spiking convolutional neural network using integrate-and-fire neurons, and a spiking recurrent neural network using current-based leaky integrate-and-fire neurons. We found that NIR faithfully represents the idealized underlying computation and that the execution aligns well across the supported platforms for feed-forward dynamics, such as a single LIF neuron or a spiking convolutional network. For more complex and time-sensitive models, such as the recurrent network, we found that platform-specific constraints and discretization choices produced somewhat diverging activation patterns and accuracies. Although the graphs are compatible, the continuous-time nature of the primitives requires platform-dependent optimization, such as quantization-aware training, on digital systems to achieve the highest accuracies. The current version of NIR excludes other important mechanisms such as adaptive threshold mechanisms [75], gating [76, 77], resonate-and-fire [78, 66] and multicompartmental [79, 80] neuron models. That said, the open design of NIR allows for easy integration of new primitives, ensuring that NIR can adapt to new applications based on neuronal computing without changing other system-level aspects.
In the context of neuromorphic computing, the notion of mismatch has traditionally been associated with the imperfections and variability inherent in analog hardware systems. However, our findings highlight a crucial and often overlooked aspect of mismatch in purely digital neuromorphic systems, particularly when working with recurrent neural networks. This underscores the need for strategies that address mismatch across both analog and digital neuromorphic systems to ensure consistent computational outcomes and enhance system reliability across various applications.
One key aspect that we do not discuss in this work is the connection to physical systems and future hardware platforms, including analog. NIR can represent models of arbitrary scale, limited only by the underlying hardware of (1) the system that converts a model to and from the NIR format, and (2) the neuromorphic hardware that runs such a model. Both the SpiNNaker2 and Loihi platforms scale to multi-chip systems, which is necessary to address larger problems such as energy-efficient large models [81]. For example, Hala Point combines 1.15 billion neurons distributed across over 1k Loihi 2 processors. Additionally, NIR can compose into highly complex systems of equations, such as higher-order systems, which lends itself well to the study of diffusion, fluid dynamics, and optimization problems for a given architecture-hardware-software system. NIR primitives can be realized directly as physical circuit components and can therefore generalize to a richer set of hardware systems than the ones we have discussed so far. Mapping such models onto larger and richer hardware platforms would be exciting to explore in future work.
It is crucial that the computations represented by NIR graphs remain the same, despite evolving primitives and future developments. An important step towards this is the continued testing of platform integrations along with a well-documented versioning approach for the primitives akin to the Operator Set (opset) number in ONNX [14]. Additionally, and specifically for digital hardware, it would be interesting to explore the integration of lower-level hardware details for multi-level models that further optimize efficiency, similar to MLIR dialects [15].
The current computational experiments were chosen to demonstrate the capabilities of NIR as proof-of-concept of the representation’s modular functionality. Hence, the tasks such as N-MNIST with SCNN and Braille dataset with SRNN are not challenging for current ML algorithms, and we prioritized these illustrations on state-of-the-art neural network architectures. Yet, the primitives used in the models are highly relevant for neuromorphic computing [82], since recent work on CNNs and ResNets is successfully competing with ANNs for vision [83, 84, 85], and for time-series classification [86]. These SNNs use far less computing than their ANN counterparts, which can significantly reduce energy consumption when implemented on dedicated hardware [87]. Large-scale neuromorphic systems like Loihi or SpiNNaker2 and others have shown orders of magnitude energy and latency improvements compared to off-the-shelf CPUs and GPUs, see Refs. [88, 89] for reviews. As the deep learning community has moved to transformers as a primary component, also efficient, brain-inspired architectures such as SpikeGPT [90] have been developed. There may be specific applications that run on these current architectures that NIR may add complex operations, which may reduce energy efficiencies. For these systems, we do not recommend using NIR as means to increase efficiencies. However, for many applications, NIR can become a useful tool for cross-platform interoperability by providing flexibility in choosing the most efficient hardware platform for any given neural architecture. In addition, depending on the application, NIR may reduce the number of training operations with a potential increase in energy efficiency.
Our main goal has been to address the lack of shared representations for neuromorphic computing. The deployment of software-generated models on neuromorphic hardware was previously only possible through hardware-specific software libraries. NIR effectively decouples the development of neural models from platform-specific tools, such that software and hardware can evolve independently, as shown in Figure 2, and such that models developed in one framework can be readily deployed and reproduced across neuromorphic simulators and hardware within device-specific boundary conditions. In turn, NIR allows users to directly leverage tools and resources from the active and growing neuromorphic open-source community, such as quantization mechanisms or algorithms for fast training. Furthermore, interoperability of spiking neuron models simplifies benchmarking and comparison across platforms. Any framework that chooses to integrate with an intermediate representation, such as NIR, benefits from these tools as well as a much larger user base compared to standalone technology stacks. The present set of primitives is limited, and many platforms are not included in our analysis. However, similar to how the initial design of digital instruction sets propelled the advent of modern-day computing, we believe this to be a crucial first step towards further development of shared tooling infrastructure for neuromorphic computing. We have argued that abstracting away hardware-specific details will lower the barrier to entry for new neuromorphic hardware systems, and it is our hope that NIR can accelerate the academic development required to propel neuromorphic computing to perform on-par with present-day deep learning, while solving practical problems in the near future.